# access point PTP Precision Time Protocol

### v1

@@ VII Results
We build a RAN network with $M=8$ cells with $N=99$ UEs in each cell. Thus, we have eight twins in the TwinRAN-multi instance and 100 twins in each of the 8 TwinRAN-cell instances. Each cell in the TwinRAN-multi instance is updated five times a second, while each Twin in the TwinRAN-cell instance is updated 10 times a second. For each of the three use cases, we define relevant KPIs that we measure. For the first use case, the volume of data that traverses a data history connection per hour and is stored in the database is measured. The lag from the sourceTime to the TwinRAN updateTime is measured for use case II. Finally, the time it takes to spawn five new copies of the TwinRAN instance is measured for the last use case. We also present a cost analysis taking the pricing information by Azure for the Eastern United States for all three applications.
The Azure DT generates 30 million messages and 4 million operations per hour in operation mode. As of the writing of this paper, Azure charges 1\$ per million messages and 2.5\$ per million operations, giving the base running price of our network as 40\$ per hour.
The xApp generates 240kB of data, and the event hub generates 64,000 events per second on average for the first use case. The data explorer caches each event of  21 bytes and batch-ingested into the database once every 5 minutes. Thus, approx. 4MB of data is ingested every 5 minutes. For such small amounts of data, a single event hub instance and 8vCPUs on the data explorer cluster are sufficient, as the CPU and cache utilization never exceeded 20 percent. The storage space of 1.5 TB will be saturated after three and half years of operation. An event hub with two throughput units charges 0.06\$/hour and 0.03\$/million messages, incurring an hourly charge of approx. 7\$ for the The hourly rate for the data explorer for the given specifications is 1.5\$ for the pay as you go tier. The first use case requires 48.5\$/hour, with Azure DT taking up most of the cost.
The lag of the location parameter in TwinRAN was measured for the second use case using the method explained in SectionIV. The update from the source first acts on its twin then traverses the event hub and is routed to TwinRAN. With 10,000 measurements, the average lag for this setup was found to be 140 ms, with the maximum being 562 ms. If we update the location of the UE 5 times a second, this lag should still allow the latest location information at the TwinRAN instance. However, an object moving at 200 km/hr would have moved roughly 10 meters in 200 ms. Therefore, the high latency of the Twin-Twin update would restrict the second use case to slowly moving objects such as AGVs. The cost of running the second use case is 47\$/hour - the same as the first minus the 1.5\$ for the data explorer cluster.
@@
VIII Conclusion
In this paper, we presented a concept and architecture of twinning the 5G RAN on the cloud using Azure DT. The architecture focuses on achieving three main properties of TwinRAN: vendor-agnostic, multipurpose, and non-invasive. TwinRAN incorporates and utilizes the O-RAN concept, seamlessly integrating it into a 5G RAN. Thorough investigations of the behavior of Azure DTs help us make design decisions for TwinRAN, such as maintaining two different instances simultaneously for intercell and intracell operations. The investigations also show some limitations - it is not a suitable solution for tightly coupled close-loop operations requiring latencies smaller than 100 ms. We demonstrate three use cases where TwinRAN is be used, showing its multipurpose nature. We analyze the costs of running TwinRAN under the currently openly available cloud pricing terms. Analysis shows that the operating costs are dominated by the prices of maintaining the DT in the cloud at 0.05\$ per user per hour while the use cases themselves do not add significantly to it. Our contribution underscores the feasibility and advantages of DT technology in the context of 5G networks. It sets the stage for future network management and optimization innovations using cloud-based digital twins. Our architecture aims to facilitate more intelligent, resilient, and adaptive 5G networks by bridging the gap between physical and digital network environments.

### v2

@@ VII Results
We build a RAN network with $M=8$ cells with $N=99$ UEs in each cell. Thus, we have eight twins in the TwinRAN-multi instance and 100 twins in each of the 8 TwinRAN-cell instances. Each cell in the TwinRAN-multi instance is updated five times a second, while each Twin in the TwinRAN-cell instance is updated 10 times a second. For each of the three use cases, we define relevant KPIs that we measure. For the first use case, the volume of data that traverses a data history connection per hour and is stored in the database is measured. The lag from the sourceTime to the TwinRAN updateTime is measured for use case II. Finally, the time it takes to spawn five new copies of the TwinRAN instance is measured for the last use case. We also present a cost analysis taking the pricing information by Azure for the Eastern United States for all three applications. The Azure DT generates 30 million messages and 4 million operations per hour in operation mode. As of the writing of this paper, Azure charges 1\$ per million messages and 2.5\$ per million operations, giving the base running price of our network as 40\$ per hour.
For the first use case, the xApp generates 240kB of data, and the event hub generates 64,000 events per second on average for the first use case. This was measured by monitoring the output port of the xApp and the logs from Azure event hub. The data explorer caches each event of  21 bytes and batch-ingested into the database once every 5 minutes. Thus, approx. 4MB of data is ingested every 5 minutes. For such small amounts of data, a single event hub instance and 8vCPUs on the data explorer cluster are sufficient, as the CPU and cache utilization never exceeded 20 percent throughout our experiment. The storage space of 1.5 TB will be saturated after three and half years of operation. An event hub with two throughput units charges 0.06\$/hour and 0.03\$/million messages, incurring an hourly charge of approx. 7\$ for the The hourly rate for the data explorer for the given specifications is 1.5\$ for the pay as you go tier. The first use case requires 48.5\$/hour, with Azure DT taking up most of the cost.
The lag of the location parameter in TwinRAN was measured for the second use case using the method explained in SectionIV. The update from the source first acts on its twin, then traverses the event hub and is routed to TwinRAN. With 10,000 measurements, the average lag for this setup was found to be 140 ms, with the maximum being 562 ms. If we update the location of the UE 5 times a second, this lag should still allow the latest location information at the TwinRAN instance. However, an object moving at 200 km/hr would have moved roughly 10 meters in 200 ms. Therefore, the high latency of the Twin-Twin update would restrict the second use case to slowly moving objects such as AGVs. The cost of running the second use case is 47\$/hour - the same as the first minus the 1.5\$ for the data explorer cluster.
@@
Use Case KPI Name KPI Value Cost of Running
1 Data Volume 48MB/hour 48.5\$/hour
2 Twin-Twin Update Latency 140 ms 47\$/hour
3 Spawning Time for 5 copies 22 s 40\$/hour/copy
TABLE II: Results showing each use case’s selected KPI and pricing. Most of the cost comes from running the Azure DT at full capacity in the cloud at 40\$ per hour. Adding applications (use cases) is relatively cheaper, as seen from the first and second rows. Hence, scaling TwinRAN to more use cases can increase cost efficiency. This scaling is easier due to the non-invasive, multipurpose, and vendor-agnostic characteristics of TwinRAN.
VIII Limitations and Conclusion
In this paper, we presented a concept and architecture for twinning the 5G RAN on the cloud using Azure DT. The architecture focuses on achieving three main properties of TwinRAN: vendor-agnostic, multipurpose, and non-invasive. TwinRAN incorporates and utilizes the O-RAN concept, seamlessly integrating it into a 5G RAN. Thorough investigations of the behavior of Azure DTs help us make design decisions for TwinRAN, such as maintaining two different instances simultaneously for intercell and intracell operations. The investigations also show some limitations - it is not a suitable solution for tightly coupled close-loop data-driven operations requiring latencies smaller than 100 ms, such as ChARM[62] for spectrum sharing. We demonstrate three use cases where TwinRAN is used, showing its multipurpose nature. We analyze the costs of running TwinRAN under the currently openly available cloud pricing terms. Analysis shows that the operating costs are dominated by the prices of maintaining the DT in the cloud at 0.05\$ per user per hour while the use cases themselves do not add significantly to it. These costs currently make it untenable for operators to deploy TwinRAN all the time over their entire network. Therefore, with the current pricing structure, TwinRAN might be beneficial to be deployed for a partial part of an operator’s network for a limited amount of time to tackle specific problems like heavy traffic, what-if studies, or collecting data for training or publishing. The same TwinRAN instance can execute all three use cases simultaneously, and costs do not increase as more applications are deployed. We believe that if an operator were to implement TwinRAN, they could negotiate with Azure for a subscription at better pricing and scale it to many more instances. In this work, we are limited to 100 twins/instance due to the update rate, as mentioned in Section V-A, and to 8 instances due to the tier-level subscription to Azure DT provided by our academic institution.
Our contribution underscores the feasibility and advantages of DT technology in the context of 5G networks. It sets the stage for future network management and optimization innovations using cloud-based digital twins. Our architecture aims to facilitate more intelligent, resilient, and adaptive 5G networks by bridging the gap between physical and digital network environments.